OpenStudy (anonymous):

Hi guys, Help me to find q(x)? X3+2x2-29x-3 = (x+1)q(x)

OpenStudy (anonymous):

\[q(-1) = (-1)^3 + 2(-1)^2 - 29(-1) \] \[\\\\\ = -1 + 2 - 29 \]

OpenStudy (anonymous):

SO, x+1 is not a factor

OpenStudy (anonymous):



OpenStudy (anonymous):

Okay got you question!

OpenStudy (anonymous):

sorry typo error in last qns!!! this one correct.. X3+2x2-29x-30 = (X+1)(x2+x-30) q(x)=(x2+x-30) ? I want to know how they make x2+x-30?

OpenStudy (anonymous):

http://www.twiddla.com/841565

OpenStudy (anonymous):

go there its solved there

OpenStudy (anonymous):

I connect through mobile, so it take too much time to open this site.. For what that site yar?

OpenStudy (anonymous):

YAR means

OpenStudy (anonymous):

friend.

OpenStudy (anonymous):

Got that

OpenStudy (anonymous):

which COUNTRY u belong to

OpenStudy (anonymous):

India. U?

OpenStudy (anonymous):

yes I got,,but I can't understand

OpenStudy (anonymous):

England

OpenStudy (anonymous):

what u studying?

OpenStudy (amistre64):

polynomial division

OpenStudy (anonymous):

Wel come AMI

OpenStudy (amistre64):

howdy

OpenStudy (anonymous):

@amistre64 have a look http://openstudy.com/study?login#/updates/503a002fe4b0edee4f0d88db its by @TheViper

OpenStudy (amistre64):

is it another "tutorial"?

OpenStudy (anonymous):

GOOD ONE

OpenStudy (theviper):

oh thanx again :)

OpenStudy (anonymous):

Its srly Intresting

OpenStudy (amistre64):

nrop, its easier to communicate thru the posting, rather than thry the messaging

OpenStudy (anonymous):

LOL

OpenStudy (anonymous):

Messaging - PRIVATE POSTING - PUBLIC

OpenStudy (amistre64):

if you can do longhand division with numbers; you can do it with polynomials as well; its the exact same steps

OpenStudy (theviper):

I message them I m biggest fan of like u @Rohangrr :)

OpenStudy (anonymous):

Thanks for that compliment

OpenStudy (theviper):

I also messaged @amistre64 :D